# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 9.1. Comparison with constant coefficients and sufficient monotonicity
    We conclude this subsection by noting that the results presented here are part of a more general program, which exploits a general method for determining the validity of the (ZMP) for dual cone subharmonics.
@@
    The comparison principles of Theorems 9.7 and 9.13 are also corollaries of the general comparison principle for constant coefficient potential theories contained in [8, Theorem 7.5]. This result states that comparison for a constant coefficient subequation $\mathcal{F}$ on a domain $\Omega$ holds if $\mathcal{F}$ is $\mathcal{M}$-monotone for a monotonicity cone subequation $\mathcal{M}$ for which there exists a $C^{2}$-strictly $\mathcal{M}$-subharmonic function $\psi$ on $\Omega$, as described in Remark 8.10 above. The point is that the existence of $\psi$ ensures the validity of the (ZMP) for the $\widetilde{\mathcal{M}}$-subharmonics on $\Omega$. Theorems 9.7 and 9.13 correspond to the monotonicity cone subequations
@@
    respectively. These two cones do admit a (quadratic) $C^{2}$-strictly $\mathcal{M}$-subharmonic function $\psi$ on every bounded domain $\Omega$, as shown in Theorem 6.6 of [8]. Moreover, the two cones in (9.8) are two members of a fundamental family of monotonicity cone subequations which are generated by double and triple intersections of $\mathcal{M}(\mathcal{P})$ above, $\mathcal{M}(\mathcal{N}):=\mathcal{N}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ and three one-parameter cones:
    $\mathcal{M}(\gamma):=\{(r,p,A)\in\mathcal{J}^{2}:\ r\leqslant-\gamma|p|\},\quad\gamma\in(0,+\infty);$
@@
    $\mathcal{M}(R):=\left\{(r,p,A)\in\mathcal{J}^{2}:\ A\geqslant\frac{|p|}{R}I\right\},\quad R\in(0,+\infty);$
    $\mathcal{M}(\mathcal{D}):={\mathbb{R}}\times\mathcal{D}\times\mathcal{S}(n),\quad\mathcal{D}\subset{\mathbb{R}}^{n}\ \text{is a {\em directional cone}};$
@@
    that is, $\mathcal{D}$ is a closed convex cone with vertex at the origin and non-empty interior. Notice that $\mathcal{M}(\mathcal{P})$ and $\mathcal{M}(\mathcal{N})$ are the “limits” of $\mathcal{M}(R)$ and $\mathcal{M}(\gamma)$ for $R\nearrow+\infty$ and $\gamma\searrow 0$, respectively. For all of the cones in the fundamental family, the validity of the (ZMP) for the $\widetilde{\mathcal{M}}$-subharmonics on $\Omega$ and hence the vadidity of comparison for any $\mathcal{M}$-monotone potential theory is known from the detailed analysis in [8]. There is a simple dichotomy: comparison holds on every bounded $\Omega$ except for the case when the cone has a generator $\mathcal{M}(R)$ with $R$ finite, and in this case comparison holds on all domains $\Omega$ which are contained in a translate of the truncated cone $\mathcal{D}_{R}=\mathcal{D}\cap B_{R}(0)$. The family above is fundamental in the sense that any monotonicity cone subequation contains one of the fundamental cones, and hence for every constant coefficient $\mathcal{M}$-monotone potential theory, the comparison principle holds (with possibly a restriction on the size of the domain $\Omega$).
@@
    As a final consideration, the analysis discussed in Remark 9.14 has two additional extensions. First, by exploiting the correspondence principle of Theorem 11.13 of [8] one deduces the validity of the comparison principle for every constant coefficient and $\mathcal{M}$-monotone differential operator $F$ which is compatible with an $\mathcal{M}$-monotone constant coefficient subequation $\mathcal{F}$. Many applications of this are discussed in Part IV of [8]. Second, in the variable coefficient case, if $\mathcal{F}$ is also fiberegular, then the analysis above extends to the potential theory defined by $\mathcal{F}$ as well as to associated PDE for any operator $F$ which is compatible with $\mathcal{F}$. This program was carried out in [11], which also gives characterizations of $\widetilde{\mathcal{M}}$-subharmonics for the other monotonicty cones $\mathcal{M}$ in the fundamental family, extending the characterizations above for the cones $\mathcal{M}(\mathcal{P})$ and $\mathcal{M}(\mathcal{N},\mathcal{P})$ (see Section 6 of [11]) The following subsection treats some complementary results to [8] and [11].
    9.2. Strict comparison with minimal monotonicity and quasi-convex functions
@@
    In this subsection, we present some comparison principles which are not included in the aforementioned [8] and [11]. These results give additional situations in which one has sufficient monotonicity in order to prove a form of comparison. In fact, the minimal monotonicity present in every subequation becomes sufficient, if one weakens comparison to a notion of strict comparison. In the variable coefficient setting, this form of comparison is shown to hold for quasi-convex functions, while, in the constant coefficient setting, it is shown to hold for functions which are merely semicontinuous. Two approaches for the semicontinuous version will be presented, both of which use quasi-convex approximation; one in an implicit sense and one in an explicit sense, which will be compared in Remark 9.21
@@
    We begin by noting that, by properties (P) and (N), every (variable coefficent) subequation $\mathcal{F}$ has a constant coefficient monotonicity cone; namely the minimal monotonicity cone
@@
    which, a priori, might be the maximal monotonicity cone for $\mathcal{F}$. As noted previously, $\mathcal{M}_{0}$ is not a subequation because it has empty interior and hence does not satisfy the topological property (T). In such a situation, the methods described in the previous subsection do not apply. However, one can ask under what situations does $\mathcal{M}_{0}$-monotonicity for $\mathcal{F}$ still yield comparison in the potential theory determined by $\mathcal{F}$? One way is to focus on quasi-convex subharmonics and to add a strictness assumption. To that end, we give the following notion of strong strictness.
@@
    We say that $u\in\operatorname{USC}(X)$ is strongly strictly $\mathcal{F}$-subharmonic on $X$ if $u\in\mathcal{G}(X)$ for some subequation $\mathcal{G}\subset\operatorname{int}{\mathcal{F}}$. We will write $u\in\mathcal{F}_{\rm strict}(X)$.
@@
    For $n=1$ and $X=(0,1)$, consider $F(x,p)=x-p$. Then $u(x)=\frac{x^{3}}{3}$ is a strict smooth subsolution on $X$, according to Definition 5.10, but it is not strongly strict in the viscosity sense of Definition 9.15.
    Harvey and Lawson [26] showed that one always has the (ZMP) form of comparison if the subsolutions are quasi-convex and one of them is strongly strict, just like the Definitional Comparison Lemma 7.2 tells that (ZMP) holds, even without the quasi-convexity assumption, if one of the subsolutions is strict and smooth.
    Theorem 9.17 (Strict comparison for quasi-convex functions).
    Let $\Omega\Subset{\mathbb{R}}^{n}$ be a bounded domain and let $\mathcal{F}$ be a subequation on $\Omega$. Then strict comparison holds for quasi-convex functions:
    $u+v\leqslant 0\ \text{on $\partial\Omega$}\implies u+v\leqslant 0\ \text{on $\Omega$}.$
    if $u\in\operatorname{USC}(\overline{\Omega})\cap\mathcal{F}_{\rm strict}(\Omega)$ and $v\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{F}}(\Omega)$ are locally quasi-convex on $\Omega$.
@@
    Suppose, by contradiction, that strict comparison fails on $\Omega\Subset{\mathbb{R}}^{n}$. Then $w=u+v$ must have a positive interior maximum $m\vcentcolon=w(x_{0})>0$ at $x_{0}\in\Omega$. Hence $w$ has $(0,0)$ as an upper contact jet at $x_{0}$ and the Theorem 2.16 on summands yields
    $(u(x_{0}),p,A)\in\mathcal{G}_{x_{0}}\qquad\text{and}\qquad(v(x_{0}),q,B)\in\widetilde{\mathcal{F}}_{x_{0}},$
@@
    Theorem 9.17 is indeed useful if we can assure a priori that subsolutions and dual subsolutions of a given equation are quasi-convex. For example, consider the Monge–Ampère equation
@@
    for some $f\in\operatorname{LSC}(\overline{\Omega})$ which is positive and bounded from above. If we look for convex solutions, we immediately see that they are strongly convex, in the sense that the minimum eigenvalue $\lambda_{1}({D^{2}u})\geqslant C$ on $\Omega$, for some $C>0$; indeed by semicontinuity and compactness we know that $f\geqslant c$ on $\Omega$, for some $c>0$, while $\det D^{2}u\leqslant M$ for some upper bound $M>0$ on $f$. Therefore, if we restrict ourselves to strongly convex sub– and supersolutions, we trivially have that subsolutions are quasi-convex. Also, since the fibers of the subequation associated to (9.9) are
@@
    hence the dual subsolutions we are interested in are $\Lambda$-quasi-convex.
    The assumption that $u$ and $v$ are quasi-convex can be dropped if $\mathcal{F}$ and $\mathcal{G}$ are constant coefficient subequations. The following result is Corollary C.3 of [24].
@@
    Let $\mathcal{F}$ be a constant coefficient subequation. Then the strict comparison holds on every bounded domain $\Omega\Subset{\mathbb{R}}^{n}$:
    if $u\in\operatorname{USC}(\overline{\Omega})\cap\mathcal{F}_{\rm strict}(\Omega)$ and $v\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{F}}(\Omega)$.
    We will omit the short proof, which one finds in [24], noting only that it follows the same argument of its quasi-convex counterpart (Theorem 9.17), where the Theorem in Summands 2.16 for (locally) quasi-convex functions is replaced by a potential-theoretic version of the Theorem on Sums 4.12 for semicontinuous functions. Eventhough we will sketch an alternate proof of Theorem 9.19, for the benefit of the reader and in order to comment on the two approaches, we present Harvey and Lawson’s potential-theoretic version (which is Theorem C.1 of [24]).
    Theorem 9.20 (On Sums, simplified potential theoretic version).
@@
    $u(x_{\varepsilon})+v(y_{\varepsilon})\searrow u(x_{0})+v(x_{0}),$ (9.11)
    $|x_{\varepsilon}-y_{\varepsilon}|=o(\sqrt{\varepsilon}\,)$ (9.12)
@@
    Let $w(x,y)\vcentcolon=u(x)+v(y)$. For any $\varepsilon>0$, let $(x_{\varepsilon},y_{\varepsilon})$ be a maximum of $w(x,y)-Q_{\varepsilon^{-1}I}(x-y)$ on $\overline{\Omega}\times\overline{\Omega}$; by [13, Proposition 3.7], by taking $\varepsilon$ small enough, we may suppose that $(x_{\varepsilon},y_{\varepsilon})\in\Omega\times\Omega$. Then by Theorem 4.12 there exist $A_{\varepsilon},B_{\varepsilon}$ such that (9.10) holds since $u\in\mathcal{F}(X)$, $v\in\mathcal{G}(X)$ and the fibers of subequations are closed, and
@@
    $-\frac{1}{\varepsilon}\bigl{(}1+\|J\|\bigr{)}\,I\leqslant\left(\!\begin{array}[]{cc}A_{\varepsilon}&0\\ 0&B_{\varepsilon}\\ \end{array}\!\right)\leqslant\frac{1}{\varepsilon}\bigl{(}J+J^{2}\bigr{)},$
@@
    which is easily seen to be (9.19). Finally, all the remaining properties to prove easily follow again from [13, Proposition 3.7]. ∎
@@
    We conclude with some reflections on the role of quasi-convex approximation in strict comparison under minimal monotonicity.
    Remark 9.21 (Strict comparison and quasi-convex approximation).
@@
    The proof of Theorem 9.19 above is an “almost conventional” viscosity theory proof in that it exploits a potential-theoretic formulation of the conventional Theorem on Sums (based on the doubling of variables and penalization). Quasi-convex approximation (by way of the supconvolution) appears implicitly in the proof, since it plays a crucial role in the proof of the Theorem on Sums.
@@
    One might ask if an explicit use of the quasi-convex approximation technique can reduce Theorem 9.19 to the quasi-convex version of Theorem 9.17. If so, one would be extending the quasi-convex approximation technique from sufficient monotonicity $\mathcal{M}$ (as discussed in the previous subsection) to minimal monotonicity $\mathcal{M}_{0}$, at least for strict comparison. One encounters a possibile obstruction to this intent; namely, one would like to know that the quasi-convex approximations preserve subharmonicity for the subequations $\mathcal{G}$ and $\widetilde{\mathcal{F}}$. Proposition 8.8 says that this is true, provided that such subharmonics are bounded.
    Boundedness from above on compact sets is ensured by the upper semicontinuity of the subharmonics. The problem is that they may not be bounded from below. The decreasing limits property of $\mathcal{F}$-subharmonics in Proposition 7.4 v) suggests that one might first approximate an $\mathcal{F}$-subharmonic $u$ with a deceasing sequence $\{u_{m}\}_{m\in N}$ of functions which are bounded from below and $\mathcal{F}$-subharmonic. Simple truncation from below with $u_{m}:=\max\{u,-m\}$ works if each constant $m\in N$ is an $\mathcal{F}$-subharmonic by the maximum property of Proposition 7.4 ii). More generally, as seen in the proof proof of Corollary 8.9, one can take $u_{m}:=\max\{u,\varphi-m\}$ for any fixed $\mathcal{F}$-subharmonic which is bounded from below. Unfortunately, the existence of such a bounded subharmonic $\varphi$ is, in general, only guaranteed locally (as discussed in Remark 5.11).
@@
    Therefore, in a proof of strict comparison with minimal monotonicity $\mathcal{M}_{0}$ by an explicit quasi-convex approximation, one should first show that one can assume without loss of generality that the subharmonics are bounded (for example, by proving the existence of a bounded global subharmonic for $\mathcal{G}$ and $\widetilde{\mathcal{F}}$) or find a way to carry the global information of the behavior of $u+v$ on $\partial\Omega$ into a neighborhood of some given point (for instance, a point in $\Omega$ at which $u+v$ assumes its positive maximum, if one is proceeding by contradiction). If instead, one tries to work only locally as in the proof of Theorem 9.17, then the most immediate path probably leads to a proof with the flavour of the Theorem on Sums, since it is likely that one eventually ends up needing results like Theorem 2.16 and Proposition 4.11.
@@
    By contradiction, suppose that strict comparison fails for some $\Omega\Subset{\mathbb{R}}^{n}$; hence there exists $\overline{x}\in\Omega$ such that $u(\overline{x})+v(\overline{x}):=M>0$. For each $\eta\in(0,M)$ fixed, by the sliding property of Proposition 7.4 iv), $\hat{u}\vcentcolon=u-\eta\in\mathcal{G}(\Omega)$ and one still has $\hat{u}(\overline{x})+v(\overline{x})>0$. Now, with $\Phi$ and $\Psi$ satisfying (9.20), define for each for $m\in\mathbb{N}$
@@
    Note that $u_{m}$ and $v_{m}$ are bounded on $\overline{\Omega}$, $u_{m}\searrow u$ and $v_{m}\searrow v$ pointwise on $\overline{\Omega}$ as $m\to+\infty$, and $u_{m}(\overline{x})+v_{m}(\overline{x})\geqslant\hat{u}(\overline{x})+v(\overline{x})>0$ for all $m\in\mathbb{N}$. In addition, one has $u_{m}\in\mathcal{G}(\Omega)$ and $v_{m}\in\widetilde{\mathcal{F}}(\Omega)$, by the maximum property. Quasi-convex approximation (Proposition 8.8) yields the supconvolutions $u_{m}^{\varepsilon}\in\mathcal{G}(\Omega_{\delta})$ and $v_{m}^{\varepsilon}\in\widetilde{\mathcal{F}}(\Omega_{\delta})$, where
    $\delta=\delta(\varepsilon)\vcentcolon=2\sqrt{\varepsilon\max\{\max_{\overline{\Omega}}|u_{m}|,\max_{\overline{\Omega}}|v_{m}|\}}.$
    These quasi-convex approximations satisfy $u_{m}^{\varepsilon}(\overline{x})+v_{m}^{\varepsilon}(\overline{x})\geqslant u_{m}(\overline{x})+v_{m}(\overline{x})>0$ for each small $\varepsilon>0$. Therefore, by the strict comparison for quasi-convex functions (Theorem 9.17), for any $\varepsilon>0$ so small that $\overline{x}\in\Omega_{\delta}$, there exists a point $y_{\varepsilon}$ such that

### v3

@@ 9.2. Comparison with constant coefficients and sufficient monotonicity
We conclude this section by noting that the results presented here are part of a wider program, which exploits a general method for determining the validity of the zero maximum principle for dual cone subharmonics.
@@
The comparison principles of Theorems 9.8 and 9.15 are also corollaries of the general comparison principle for constant coefficient potential theories contained in [9, Theorem 7.5]. This result states that comparison for a constant coefficient subequation $\mathcal{F}$ on a domain $\Omega$ holds if $\mathcal{F}$ is $\mathcal{M}$-monotone for a monotonicity cone subequation $\mathcal{M}$ for which there exists a $C^{2}$-strictly $\mathcal{M}$-subharmonic function $\psi$ on $\Omega$, as described in Remark 8.10 above. The point is that the existence of $\psi$ ensures the validity of (ZMP) for the $\widetilde{\mathcal{M}}$-subharmonics on $\Omega$.
With the notation of [9], Theorems 9.8 and 9.15 correspond to the monotonicity cone subequations
@@
respectively. These two cones do admit a (quadratic) $C^{2}$-strictly $\mathcal{M}$-subharmonic function $\psi$ on every bounded domain $\Omega$, as shown in [9, Theorem 6.6]. Moreover, the two cones in (9.5) are two members of a fundamental family of monotonicity cone subequations which are generated by double and triple intersections of $\mathcal{M}(\mathcal{P})$ above, $\mathcal{M}(\mathcal{N})\vcentcolon=\mathcal{N}\times{\mathbb{R}}^{n}\times\mathcal{S}(n)$ and the three one-parameter cones
@@
$\begin{gathered}\mathcal{M}(\gamma)\vcentcolon=\{(r,p,A)\in\mathcal{J}^{2}:\ r\leqslant-\gamma|p|\},\quad\gamma\in(0,+\infty),\\ \mathcal{M}(R)\vcentcolon=\left\{(r,p,A)\in\mathcal{J}^{2}:\ A\geqslant\frac{|p|}{R}I\right\},\quad R\in(0,+\infty),\\ \mathcal{M}(\mathcal{D})\vcentcolon={\mathbb{R}}\times\mathcal{D}\times\mathcal{S}(n),\quad\mathcal{D}\subset{\mathbb{R}}^{n}\ \text{is a {\em directionalcone}},\end{gathered}$
a directional cone being a closed convex cone with vertex at the origin and non-empty interior. Notice that $\mathcal{M}(\mathcal{P})$ and $\mathcal{M}(\mathcal{N})$ are the “limits” of $\mathcal{M}(R)$ and $\mathcal{M}(\gamma)$ for $R\nearrow+\infty$ and $\gamma\searrow 0$, respectively.
@@
For all of the cones in the fundamental family, the validity of the zero maximum principle for the $\widetilde{\mathcal{M}}$-subharmonics on $\Omega$, and hence the validity of comparison for any $\mathcal{M}$-monotone potential theory, is known from the detailed analysis in [9]. There is a simple dichotomy: comparison holds on every bounded $\Omega$ except for the case when the cone has a generator $\mathcal{M}(R)$ with $R$ finite, and in this case comparison holds on all domains $\Omega$ which are contained in a translate of the truncated cone $\mathcal{D}\cap B_{R}(0)$.
The family above is fundamental in the sense that any monotonicity cone subequation contains one of the fundamental cones, and hence for every constant coefficient $\mathcal{M}$-monotone potential theory, the comparison principle holds (with possibly a restriction on the size of the domain $\Omega$).
@@
As a final consideration, the analysis discussed in Remark 9.16 has two additional extensions. First, by exploiting the correspondence principle of [9, Theorem 11.13] one deduces the validity of the comparison principle for every constant coefficient and $\mathcal{M}$-monotone differential operator $F$ which is compatible with an $\mathcal{M}$-monotone constant coefficient subequation $\mathcal{F}$. Many applications of this are discussed in [9, Part IV]. Second, in the variable coefficient case, if $\mathcal{F}$ is also fiberegular, then the analysis above extends to the potential theory defined by $\mathcal{F}$ as well as to associated PDE for any operator $F$ which is compatible with $\mathcal{F}$. This program was carried out in [12], which also gives characterizations of $\widetilde{\mathcal{M}}$-subharmonics for the other monotonicty cones $\mathcal{M}$ in the fundamental family, extending the characterizations above for the cones $\mathcal{M}(\mathcal{P})$ and $\mathcal{M}(\mathcal{N},\mathcal{P})$; see [12, Section 6].
The following section treats some complementary results to [9] and [12].
9.3. Strict comparison with minimal monotonicity and semiconvex functions
@@
In this section, we present some comparison principles which are not included in the aforementioned [9] and [12]. These results give additional situations in which one has sufficient monotonicity in order to prove a form of comparison. In fact, the minimal monotonicity present in every subequation becomes sufficient if one weakens comparison to a notion of strict comparison. In the variable coefficient setting, this form of comparison is shown to hold for semiconvex functions, while, in the constant coefficient setting, it is shown to hold for functions which are merely semicontinuous. Two approaches for the semicontinuous version will be presented, both of which use semiconvex approximation; one in an implicit sense and one in an explicit sense, which will be compared in Remark 9.23.
@@
We begin by recalling that, by properties (P) and (N), every (variable coefficient) subequation $\mathcal{F}$ has a constant coefficient monotonicity cone; namely the minimal monotonicity cone
@@
which, a priori, might be the maximal monotonicity cone for $\mathcal{F}$. As noted previously, $\mathcal{M}_{0}$ is not a subequation because it has empty interior and hence does not satisfy the topological property (T). In such a situation, the methods described in the previous section do not apply. However, one can ask under what situations $\mathcal{M}_{0}$-monotonicity for $\mathcal{F}$ still yields comparison in the potential theory determined by $\mathcal{F}$. One way is to focus on semiconvex subharmonics and to add a strictness assumption; to that end, we give the following notion of strong strictness.
@@
[…]
@@
For $n=1$ and $X=(0,1)$, consider $F(x,p)=x-p$. Then $u(x)=\frac{x^{3}}{3}$ is a strict smooth subsolution on $X$, according to Definition 5.10, but it is not strongly strict in the sense of Definition 9.17.
Harvey and Lawson [29] showed that one always has the form (9.1) of comparison if the subsolutions are semiconvex and one of them is strongly strict, just like the Definitional Comparison Lemma 7.3 tells that (9.1) holds, even without the semiconvexity assumption, if one of the subsolutions is strict and smooth.
Theorem 9.19 (Strict comparison for semiconvex functions).
Let $\Omega\subset{\mathbb{R}}^{n}$ be a bounded domain and let $\mathcal{F}$ be a subequation on $\Omega$. Then strict comparison holds for semiconvex functions; that is,
$\begin{dcases}u\in\operatorname{USC}(\overline{\Omega})\cap\mathcal{F}_{\rm strict}(\Omega)\cap\mathsf{sc}_{\mathrm{loc}}(\Omega)\\ v\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{F}}(\Omega)\cap\mathsf{sc}_{\mathrm{loc}}(\Omega)\\ u+v\leqslant 0\ \text{on $\partial\Omega$}\end{dcases}\quad\implies\quad u+v\leqslant 0\ \text{on $\Omega$}.$
@@
Suppose, by contradiction, that strict comparison fails on $\Omega\Subset{\mathbb{R}}^{n}$. Then $w=u+v$ must have a positive interior maximum $m\vcentcolon=w(x_{0})>0$ at $x_{0}\in\Omega$. Hence $w$ has $(0,0)$ as an upper contact jet at $x_{0}$ and Theorem 2.16 on summands yields
$(u(x_{0}),p,A)\in\mathcal{G}_{x_{0}},\qquad(v(x_{0}),q,B)\in\widetilde{\mathcal{F}}_{x_{0}},$
@@
Theorem 9.19 is indeed useful if we can assure a priori that the subharmonics and dual subharmonics of a given subequation are semiconvex. For example, consider the subequation associated to the inhomogeneous Monge–Ampère equation
@@
for some $f\in C(\Omega)$ which is positive and bounded from above. If we look for convex solutions $u$, we immediately see that they are strongly convex, in the sense that $\lambda_{1}({D^{2}u})\geqslant C$ on $\Omega$, for some $C>0$; indeed by semicontinuity and compactness we know that $f\geqslant c$ on $\Omega$, for some $c>0$, while $\det D^{2}u\leqslant M$ for some upper bound $M>0$ on $f$. Therefore, if we restrict ourselves to strongly convex sub– and supersolutions, we trivially have that subsolutions are semiconvex. Also, since the fibers of the subequation associated to (9.6) are
@@
hence the dual subsolutions we are interested in are $\Lambda$-semiconvex.
The assumption that $u$ and $v$ are semiconvex can be dropped if $\mathcal{F}$ and $\mathcal{G}$ are constant coefficient subequations. The following result is [27, Corollary C.3].
@@
Let $\mathcal{F}$ be a constant coefficient subequation. Then strict comparison holds on every bounded domain $\Omega\subset{\mathbb{R}}^{n}$; that is,
$\begin{dcases}u\in\operatorname{USC}(\overline{\Omega})\cap\mathcal{F}_{\rm strict}(\Omega)\\ v\in\operatorname{USC}(\overline{\Omega})\cap\widetilde{\mathcal{F}}(\Omega)\\ u+v\leqslant 0\ \text{on $\partial\Omega$}\end{dcases}\quad\implies\quad u+v\leqslant 0\ \text{on $\Omega$}.$
We will omit the short proof, which one finds in [27], noting only that it follows the same argument of its semiconvex counterpart (Theorem 9.19), where the Theorem in Summands 2.16 for locally semiconvex functions is replaced by a potential-theoretic version of the Theorem on Sums 4.13 for semicontinuous functions. Even though we will sketch an alternate proof of Theorem 9.21, for the benefit of the reader and in order to comment on the two approaches, we present Harvey and Lawson’s potential-theoretic version of the Theorem on Sums [27, Theorem C.1].
Theorem 9.22 (On Sums; simplified potential theoretic version).
@@
$u(x_{\varepsilon})+v(y_{\varepsilon})\searrow u(x_{0})+v(x_{0}),$
$|x_{\varepsilon}-y_{\varepsilon}|=o(\sqrt{\varepsilon}\,),$
@@
Let $w(x,y)\vcentcolon=u(x)+v(y)$. For any $\varepsilon>0$, let $(x_{\varepsilon},y_{\varepsilon})$ be a maximum of $w(x,y)-Q_{\varepsilon^{-1}I}(x-y)$ on $\overline{\Omega}\times\overline{\Omega}$; by [14, Proposition 3.7], by taking $\varepsilon$ small enough, we may suppose that $(x_{\varepsilon},y_{\varepsilon})\in\Omega\times\Omega$. Then by Theorem 4.13 there exist $A_{\varepsilon},B_{\varepsilon}$ such that (9.7) holds (since $u\in\mathcal{F}(X)$, $v\in\mathcal{G}(X)$ and the fibers of subequations are closed) and
@@
$-\frac{1}{\varepsilon}\bigl{(}1+\|J\|\bigr{)}\,I\leqslant\left(\!\begin{array}[]{cc}A_{\varepsilon}&0\\ 0&B_{\varepsilon}\\ \end{array}\!\right)\leqslant\frac{1}{\varepsilon}\bigl{(}J+J^{2}\bigr{)},\qquad J\vcentcolon=\left(\!\begin{array}[]{cc}I&-I\\ -I&I\\ \end{array}\!\right),$
@@
which is easily seen to be (9.14). Finally, all the remaining properties to prove easily follow again from [14, Proposition 3.7]. ∎
@@
We conclude with some reflections on the role of semiconvex approximation in strict comparison under minimal monotonicity.
Remark 9.23 (Strict comparison and semiconvex approximation).
@@
The proof of Theorem 9.21 above is an “almost conventional” viscosity theory proof in that it exploits a potential-theoretic formulation of the conventional Theorem on Sums, based on the doubling of variables and penalization. Semiconvex approximation, by way of the sup-convolution, appears implicitly in the proof, since it plays a crucial role in the proof of the Theorem on Sums.
@@
One might ask if an explicit use of the semiconvex approximation technique can reduce Theorem 9.21 to the semiconvex version of Theorem 9.19. If so, one would be extending the semiconvex approximation technique from sufficient monotonicity $\mathcal{M}$ (as discussed in the previous section) to minimal monotonicity $\mathcal{M}_{0}$, at least for strict comparison. We note that one encounters a possible obstruction to this intent; namely, one would like to know that the semiconvex approximations preserve subharmonicity for the subequations $\mathcal{G}$ and $\widetilde{\mathcal{F}}$, and Proposition 8.8 says that this is true, provided that such subharmonics are bounded. Boundedness from above on compact sets is ensured by the upper semicontinuity of the subharmonics; the problem is that they may not be bounded from below.
[…]
@@
Therefore, in a proof of strict comparison with minimal monotonicity $\mathcal{M}_{0}$ by an explicit semiconvex approximation, one should first show that one can assume without loss of generality that the subharmonics are bounded (for example, by proving the existence of a bounded global subharmonic for $\mathcal{G}$ and $\widetilde{\mathcal{F}}$) or find a way to carry the global information of the behavior of $u+v$ on $\partial\Omega$ into a neighborhood of some given point (for instance, a point in $\Omega$ at which $u+v$ assumes its positive maximum, if one is proceeding by contradiction). If instead, one tries to work only locally as in the proof of Theorem 9.19, then the most immediate path probably leads to a proof with the flavour of the Theorem on Sums, since it is likely that one eventually ends up needing results like Theorem 2.16 and Proposition 4.12.
@@
By contradiction, suppose that strict comparison fails for some $\Omega\Subset{\mathbb{R}}^{n}$; hence there exists $\overline{x}\in\Omega$ such that $u(\overline{x})+v(\overline{x})\vcentcolon=M>0$. For each $\eta\in(0,M)$ fixed, by the sliding property (Proposition 7.5(iv)), $\hat{u}\vcentcolon=u-\eta\in\mathcal{G}(\Omega)$ and one still has $\hat{u}(\overline{x})+v(\overline{x})>0$.
Now, with $\Phi$ and $\Psi$ satisfying (9.15), define for each for $m\in\mathbb{N}$
@@
Note that $u_{m}$ and $v_{m}$ are bounded on $\overline{\Omega}$, $u_{m}\searrow u$ and $v_{m}\searrow v$ pointwise on $\overline{\Omega}$ as $m\to+\infty$, and $u_{m}(\overline{x})+v_{m}(\overline{x})\geqslant\hat{u}(\overline{x})+v(\overline{x})>0$ for all $m\in\mathbb{N}$. In addition, one has $u_{m}\in\mathcal{G}(\Omega)$ and $v_{m}\in\widetilde{\mathcal{F}}(\Omega)$, by the maximum property (Proposition 7.5(ii)). Semiconvex approximation (Proposition 8.8) yields the sup-convolutions $u_{m}^{\varepsilon}\in\mathcal{G}(\Omega_{\delta})$ and $v_{m}^{\varepsilon}\in\widetilde{\mathcal{F}}(\Omega_{\delta})$, where
$\delta=\delta(\varepsilon)\vcentcolon=2\sqrt{\varepsilon\max\Bigl{\{}\max_{\overline{\Omega}}|u_{m}|,\,\max_{\overline{\Omega}}|v_{m}|\Bigr{\}}}\,.$
These semiconvex approximations satisfy $u_{m}^{\varepsilon}(\overline{x})+v_{m}^{\varepsilon}(\overline{x})\geqslant u_{m}(\overline{x})+v_{m}(\overline{x})>0$ for each small $\varepsilon>0$. Therefore, by the strict comparison for semiconvex functions (Theorem 9.19), for any $\varepsilon>0$ so small that $\overline{x}\in\Omega_{\delta}$, there exists a point $y_{\varepsilon}$ such that